A student dance committee is to be formed consisting of 2 boys and 4 girls. If the membership is to be chosen from 7 boys and 6 girls, how many different committees are possible?
Answer by stanbon(75887)

A student dance committee is to be formed consisting of 2 boys and 4 girls. If the membership is to be chosen from 7 boys and 6 girls, how many different committees are possible?
---
# of ways to select 2 boys = 7C2 = (7*6)/(1*2) = 21
# of ways to select 4 girls = 6C4 = 6C2 = (6*5)/(1*2) = 15
-------------
Ans:: 21*15 = 315 committees
------------
Cheers,
Stan H.
-------------